### v2

# Black hole based general relativistic limit of $f(R)$ theory of gravity

Abstract
The Galactic Center black hole environment gives us new opportunity to test deviation from General Relativity and black hole physics. In this work we analytically generate the shape of the Galactic Center black hole by using a recently developed exact stationary, axisymmetric and vacuum solution of $f(R)$ gravity theory. By using scalaron mass as a free parameter we find that the shadow shape along with displacement and asymmetry is sensitive to the scalaron mass, even after keeping the black hole spin low. We recognize scalaron mass which is compatible with Kerr like quadrupole moment and hence black hole "no-hair" theorem. The same mass scale is found to reproduce the PPN parameter ($\gamma$) constrained in the weak field limit of the solar system. Gravitational identifiers, the Kretschmann scalar ($\kappa$) and gravitational potential ($\phi$) have been used to infer scalaron masses in the regime of S-stars which are found to be consistent with the limits obtained using shadow scales. We ensure that $f(R)$ gravity scalaron has an appropriate general relativistic limit in the horizon scale of the black hole. We also identify the possibility of scale invariance of the general relativistic limit.
1 Introduction
Black hole based tests of gravitational theories inform us whether there is a remarkable deviation from General Relativity (GR) in the extreme environments of spacetime. Detection of gravitational redshift of star light and Schwarzschild periapsis shift of compact stellar orbits near the Galactic Center (GC) black hole, Sgr A* (Meyer et al. (2012); Abuter et al. (2018)), horizon scale images of Sgr A* and M87* black hole (Akiyama et al. (2019, 2022a)), constraint on graviton mass (Abbott et al. (2020)) and test of Hawking’s area theorem in gravitational wave observation (Abac et al. (2025)) unfold new windows for testing GR and alternative theories of gravity. Remarkable extrapolation of GR to black hole physics and cosmology is inspired by step by step tests of the theory encompassing various astrophysical scales (Will (2014); Lalremruati and Kalita (2021)).
Cosmologies based on GR involve initial singularity (Novello and Bergliaffa (2008); Starobinsky (1980)), mysterious component violating the strong energy condition ($\rho+3p\geq 0$) required for primordial and late time accelerated expansion (Peebles and Ratra (2003); Amendola and Tsujikawa (2010)), the cosmological constant problem (Weinberg (1989)) and illusive, hitherto unidentified Cold Dark Matter particles (Blumenthal et al. (1984); Peebles (1982); Amole et al. (2016)). These are the main reasons behind research on alternatives to GR and their testability.
Black hole based tests of gravitational physics have important prospects. It is believed that astrophysical black holes are described by the Kerr metric - the stationary, axisymmetric, vacuum solutions of Einstein’s general relativistic field equations. The higher order multipole moments $(l\geq 2)$ of the external gravitational field of a black hole are determined uniquely by the mass (M) and spin ($a$) of the black hole leading to the "no-hair" theorem $Q=-Ma^{2}$ where Q is the quadrupole moment.
It was expected earlier that orbits of very short period S-stars in the nuclear star cluster near Sgr A* black hole can be used to measure angular momentum ($a$) and quadrupole moment Q of the black hole. But there are potential complications. Detection of such short period stars is associated with perturbation from other compact objects near the black hole (Sadeghian and Will (2011)). Merritt et al. (2010) showed that testing "no-hair" theorem requires stellar orbits as compact as 0.2mpc ($\sim 41$) au.
Discovery of the black hole shadow enclosed by the bright emission rings of gravitationally lensed photons near the M87* black hole and the Sgr A* black hole (Akiyama et al. (2019, 2022a)) has given naive opportunities to test the Kerr hypothesis and the black hole "no-hair" theorem directly. It is found that size of the black hole shadow is $5M\pm 7\%$ (Psaltis et al. (2020)) irrespective of the black hole spin thereby, providing a null test of the Kerr metric hypothesis. On the other hand shape of the black hole shadow is circular except for exceptionally high spin, thanks to the relation $Q=-Ma^{2}$. Significant asymmetry of the shape for low black hole spin indicates a departure from the "no-hair" theorem, hence enabling the shape of the black hole shadow to test the general relativistic "no-hair" theorem.
New tests provide us with constraints on underlying theories of gravity. These are tests involving large gravitational potential ($\phi\sim M/r$) and large spacetime curvature ($\kappa\sim M/r^{3}$, the Kretschmann scalar). These tests inform us whether there exists potential deviation from GR in the strong field regimes. It is also possible to map strong field tests of a gravitation theory to its weak field tests. Conventional methods needed to test gravitational theories in the weak field regime of the solar system is to involve theory agnostic Parametrized Post Newtonian (PPN) metric (Will (2014)). PPN parameters are constrained by measuring light deflection near the Sun and in-plane perihelion shift of the planets (Shapiro et al. (2004); Bertotti et al. (2003); Verma et al. (2014)). These are non linearity of gravity ($\beta$) and spatial curvature produced per unit mass (the Eddington-Robertson-Schiff parameter, $\gamma$). Gravitational tests across various scales provide clue to whether a fundamental theory of gravity is scale invariant (Akiyama et al. (2022b)). For black hole based tests involving horizon scale imaging observation, one considers theory agnostic non-Kerr metrics free from pathologies (Johannsen and Psaltis (2011); Vigeland et al. (2011); Johannsen (2013); Cardoso et al. (2014); Rezzolla and Zhidenko (2014); Konoplya et al. (2016)). The parametric deformation of the Kerr metric need not be vacuum solutions of GR. One then constrains the deviations from the Kerr metric. To map strong field tests with the weak field tests, one considers PN expansion of the non-Kerr metric near black holes and takes a conservative hypothesis that weak field constraints on the PPN parameters in the solar system ($\beta-\gamma=0,\beta=1=\gamma$) are valid in the black hole environment (Psaltis et al. (2020)).
There are two direct approaches to test an underlying theory of gravity. First, the parametrically deformed Kerr metric is mapped to known black hole solutions of specific alternative gravitational theories such as Randall-Sundram II braneworld black holes (Randall and Sundrum (1999)), Einstein-Dilaton-Gauss-Bonnet black holes (Yunes and Stein (2011)) and dynamical Chern-Simons black holes (Yunes and Pretorius (2009)). This puts constraint on parameters of those theories. Second, one takes a particular theory of gravity and compares the constraints on parameters of the stationary, axisymmetric and vacuum solution with those measured in various scales. This tells us if the theory has some preferred scale. Here, we adopt the second approach.
A promising gravitational alternative to GR, which addresses late time cosmic acceleration (Capozziello and Fang (2002); Nojiri and Odintsov (2003); Carroll et al. (2004); Sotiriou and Faraoni (2010) and flat rotation curve of galaxies (Capozziello et al. (2007)) without invoking exotic dark energy and dark matter is the $f(R)$ gravity theory. Nojiri and Odintsov (2011) studied compatibility of traditional $f(R)$ and Horava-Lifshitz $f(R)$ gravity in realizing unified description of inflation and dark energy. Possibility of realizing bouncing cosmology along with unification of inflation and late time accelerated expansion of the universe in modified gravity theories has also been reviewed (Nojiri et al. (2017)). Here the gravitational field equations are obtained from the Einstein-Hilbert action with Lagrangian of the gravitational field modified as $R\longrightarrow f(R)$, $f$ being a function of the Ricci scalar. It is written as (in the unit of $G=c=1$)
$S=\frac{1}{16\pi}\int\sqrt{-g}\ d^{4}xf(R)+S_{m}(\phi_{m};g_{\mu\nu})$ (1)
Where $S_{m}$ is the action for matter fields, $\phi_{m}$ minimally coupled to spacetime metric, $g_{\mu\nu}$. GR corresponds to $f(R)=R-2\Lambda$, with $\Lambda$ being the cosmological constant.
These theories can originate from time dependent compactification of extra dimensions in string/M+ theory (Nojiri and Odintsov (2003); Nojiri et al. (2006)) and curvature corrections to quantum vacuum fluctuations (Kalita (2018, 2020)). $f(R)$ theories contain an additional scalar field $\psi=f^{\prime}(R)$ called scalaron and presents a Yukawa correction to the Newtonian potential, with range of the Yukawa force governed by the scalaron mass $M_{\psi}\sim\sqrt{1/\psi^{\prime}}$ where $\psi^{\prime}=f^{\prime\prime}(R)$ (see for example Kalita (2018)).
[…]
Hees et al. (2017) took help of 19 years of observation of short period S-stars near Sgr A* to constrain strength and scale length of Yukawa fifth force that may arise from unified theory or some models of the dark universe. Adaptive optics (AO) imaging and spectroscopic observation of two stars, S2 and S38 performed by the Keck observatory revealed that there is no deviation from GR for scales within 150 au. They showed that coupling of the fifth force to matter cannot exceed $g\approx 0.01$ at scales comparable to S2-Sgr A* distance. Recent astrometric and spectroscopic data collected through the observation of S2 has been used to constrain interaction strength of the Yukawa fifth force (Abd El Dayem et al. (2025)). The upper bound has been obtained as $|g|<0.003$ for $\lambda\sim 200$ au. It has been reported that for $\lambda<r_{S2}$ ($r_{S2}\sim$ scale of the orbit of S2) no meaningful constraint can be obtained for $g$ by considering equation of motion of the star.
It is certainly indicating a clean region of GR-a wide sub-galactic scale of its validity, without however, a clear reason for why such a theory prefers a scale above $\sim 100$ au to modify gravity. Effect of modified gravity may be hidden in horizon scales which are inaccessible to orbital studies of compact stellar orbits (see for example Paul et al. (2024)). There are other discriminators of gravitational physics. Discovery of bright emission ring near the central brightness depression observed in black hole shadow studies (M87* (Akiyama et al. (2019)) Sgr A* (Akiyama et al. (2022a))) has shown that distortion of spacetime near the black hole quantified by black hole shadow asymmetry and caused by departure from Kerr metric can be used to test GR based black hole paradigm (Psaltis et al. (2020); Johannsen and Psaltis (2010)). Theory agnostic deviation from Kerr metric is parameterized for such studies to examine impact on size and shape of black hole shadows (Johannsen and Psaltis (2011); Glampedakis and Babak (2006)). If the compact object at the GC is a black hole, and there is measurable deviation of the shape of the shadow from circularity, it indicates that perhaps we have a different theory of gravity giving rise to non-Kerr metric in the vacuum limit (Medeiros et al. (2020); Psaltis et al. (2020)). In this case it would be appropriate if one could recover GR limit of a modified theory of gravity.
Current constraints on Yukawa like fifth force in the GC environment also come from the analysis of shadow cast by the GC black hole (Vagnozzi et al. (2023); Nojiri and Odintsov (2025)) and measurement of Schwarzschild precession of S2 (Abd El Dayem et al. (2025)). The study of long range fifth force using near earth asteroids have been previously conducted (Tsai et al. (2023, 2024)). Vagnozzi et al. (2023); Khodadi et al. (2024) constrained interesting extension of GR such as string inspired spacetime, alternative theories of gravity, regular black holes, black hole mimickers, wormhole and naked singularity spacetimes through the EHT shadow of Sgr A*. The study the M87* black hole with respect to its shape and size has been studied in Bambi et al. (2019), where it was found that the Kerr bound on spin is violated.
Earlier studies on compact stellar orbits near Sgr A* (Kalita (2020); Paul et al. (2023, 2024)) and size of the shadow of Sgr A* (Kalita and Bhattacharjee (2023); Paul et al. (2024)) show that massive scalarons with $M_{\psi}>>M_{g}$, $M_{g}$ being the graviton mass ($M_{g}\sim 10^{-23}$ eV) (Abbott et al. (2021)) reproduce GR like observables. Advance of the planetary perihelia and Cassini’s bound on the PPN parameter, $\gamma$ in the solar system are found to be satisfied by scalarons which are heavier than graviton by at least 6-7 orders of magnitude (Paul and Kalita (2025)).
In this work we study black hole shadow shape by taking a new non-Kerr metric, which is stationary, vacuum and axisymmetric solution of $f(R)$ gravity theory, known as Kerr Scalaron metric developed by Paul et al. (2024). We show that $f(R)$ gravity scalaron has an appropriate GR limit while reproducing Kerr like black hole shadow measured by the Event Horizon Telescope (EHT) collaboration. We present scalaron mass which reproduces Kerr like shadow and show that scalarons respect "no-hair" theorem. It is demonstrated that scalaron mass compatible with GR like shadow matches with the ones derived from different scales such as orbits of the S-stars near Sgr A* (Kalita (2020); Paul et al. (2023)) and solar system (Paul and Kalita (2025)). These scalaron masses are found to reproduce solar system like PPN parameter, $\gamma$ in the horizon scale of the black hole. From a novel relation between scalaron mass and black hole mass, we show that scalaron mass compatible with black hole shadow (size and shape) is reproduced by curvature ($\kappa$) and potential ($\phi$) scales experienced by S-stars near the GC black hole.
The paper is organized as follows. In Section 2 we calculate the photon orbit in the Kerr Scalaron spacetime and test the "no-hair" theorem with scalarons by mapping scalaron mass onto quadrupolar correction. In Section 3 we reproduce scalaron mass from the gravitational identifiers ($\kappa$, $\phi$). We also justify the scalaron mass on the basis of an interesting observational constraint on the scalaron field amplitude. Section 4 presents the results and discussions.
2 Photon orbit with Kerr Scalaron metric and test of "no-hair theorem
2.1 The Hamilton-Jacobi formalism
The stationary, axisymmetric, vacuum metric of $f(R)$ gravity has been recently developed by employing the Newman-Janis algorithm (NJA) to the vacuum spherically symmetric and static metric of the theory Paul et al. (2024). It is known as the Kerr-Scalaron metric which has the form,
$ds^{2}=\left[1-\frac{2Mr}{\rho^{2}}\left(1+\frac{1}{3}e^{-M_{\psi}r}\right)\right]dt^{2}\\ +\frac{4Mrasin^{2}\theta}{\rho^{2}}\left(1+\frac{1}{3}e^{-M_{\psi}r}\right)dtd\phi-\frac{\rho^{2}}{\Delta}dr^{2}-\rho^{2}d\theta^{2}\\ -\left[r^{2}+a^{2}+\frac{2Mra^{2}sin^{2}\theta}{\rho^{2}}\left(1+\frac{1}{3}e^{-M_{\psi}r}\right)\right]sin^{2}\theta d\phi^{2}$ (2)
Here M is the black hole mass in the unit of $G=c=1$, where $G$ is the gravitational constant and $c$ is the speed of light. The notations $\rho$ and $\Delta$ are defined as
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta$ (3)
$\Delta=r^{2}-2Mr\left(1+1/3e^{-M_{\psi}r}\right)+a^{2}$ (4)
It was shown earlier that power law ($R^{m},m>1$) curvature correction to quantum vacuum fluctuation produces a scalaron-black hole mass relationship as $M_{\psi}\propto\frac{1}{M}$ (Kalita (2018); Talukdar et al. (2024))(see section 3). This ensures asymptotic flatness of the Kerr-scalaron metric (Paul et al. (2024)). The metric has been previously tested using pericenter shift and Lense-Thirring effect in the orbits of S-stars near GC black hole. It has been found that the heavier scalarons ($M_{\psi}\sim 10^{-17}-10^{-16}$) eV produce GR like observables Paul et al. (2024). The metric was found to likely deviate from GR for orbits smaller than 4545au. We wish to draw attention to the fact that it is possible to extract the scalaron mass from compatibility of the black hole shadow size with observed size reported by the EHT. The radius of the circular photon orbit is obtained by applying the implicit equation
$r_{ph}=2g_{tt}(r_{ph})\left(\frac{dg_{tt}}{dr}|_{r_{ph}}\right)^{-1}$ (5)
Here $g_{tt}(r)$ is the coefficient of $dt^{2}$ in the line element. The implicit equation depends on $M_{\psi}$ and the black hole spin ($a$). The diameter of the black hole shadow is obtained from the relation
$d_{sh}=\frac{2r_{ph}}{\sqrt{g_{tt}(r_{ph)}}}$ (6)
The EHT collaboration (Akiyama et al. (2022b)) measured the deviation parameter $\delta=(d_{sh})/(d_{sh,Sch})-1$ which indicates departure of the black hole shadow size from that ($6\sqrt{3}M/R_{o}$, $R_{o}$ being the distance to the black hole) expected from a Schwarzschild (Sch) black hole (general relativistic prediction). Measurement of the deviation required estimation of the mass of the Sgr A* black hole and its distance from the earth. The Very Large Telescope (VLT) and the Keck telescope independently measured the mass and distance of the black hole (Abuter et al. (2022); Do et al. (2019)). The resulting deviation parameters are $\delta=-0.08^{+0.09}_{-0.09}$ (VLT) and $\delta=-0.04^{+0.09}_{-0.10}$ (Keck). Scalaron mass $M_{\psi}=10^{-17}-10^{-16}$ eV has been found to be compatible with EHT measurements of the black hole shadow size (Paul et al. (2024)). If EHT’s observation is treated as a confirmation of the general relativistic Kerr hypothesis for black hole spacetime, it means that the above scalaron mass range is a general relativistic limit of $f(R)$ gravity theory.
In order to see the impact of the Kerr-scalaron metric on shape of the black hole shadow, we proceed by considering the Hamilton-Jacobi (HJ) approach for obtaining the null geodesics around the black hole. The HJ equation is written as
$2\frac{\partial S}{\partial\tau}=g^{ij}\frac{\partial S}{\partial x^{i}}\frac{\partial S}{\partial x^{j}}$ (7)
Here $\tau$ is the affine parameter and the $g^{ij}$ are the contravariant metric tensor components in the Kerr-scalaron metric. The HJ equation becomes,
$\begin{split}2\frac{\partial S}{\partial\tau}=\frac{\Sigma^{2}}{\Delta\rho^{2}}\left(\frac{\partial S}{\partial t}\right)^{2}+\frac{4Mar}{\Delta\rho^{2}}\left(1+\frac{1}{3}e^{-M_{\psi}r}\right)\left(\frac{\partial S}{\partial t}\right)\left(\frac{\partial S}{\partial\phi}\right)-\frac{\Delta-a^{2}\sin^{2}\theta}{\rho^{2}\Delta\sin^{2}\theta}\left(\frac{\partial S}{\partial\phi}\right)^{2}\\ \\ -\frac{\Delta}{\rho^{2}}\left(\frac{\partial S}{\partial r}\right)^{2}-\frac{1}{\rho^{2}}\left(\frac{\partial S}{\partial\theta}\right)^{2}\end{split}$ (8)
On further simplification, we obtain the following convenient form.
$\begin{split}2\frac{\partial S}{\partial\tau}=\frac{1}{\Delta\rho^{2}}\left[(r^{2}+a^{2})\left(\frac{\partial S}{\partial t}\right)+a\left(\frac{\partial S}{\partial\phi}\right)\right]^{2}-\frac{1}{\rho^{2}\sin^{2}\theta}\left[a\sin^{2}\theta\left(\frac{\partial S}{\partial t}\right)+\left(\frac{\partial S}{\partial\phi}\right)\right]^{2}\\ \\ -\frac{\Delta}{\rho^{2}}\left(\frac{\partial S}{\partial r}\right)^{2}-\frac{1}{\rho^{2}}\left(\frac{\partial S}{\partial\theta}\right)^{2}\end{split}$ (9)
Assuming that the variables are separable, we adopt a principal function of the form
$S=\frac{1}{2}\delta_{1}\tau-Et+L_{z}\phi+S_{r}(r)+S_{\theta}(\theta)$ (10)
where E and $L_{z}$ are the particle’s energy and angular momentum respectively. $S_{r}(r)$ and $S_{\theta}(\theta)$ are the radial and angular principal functions. For the chosen form of $S$, we obtain the separable equations as
$\Delta\left(\frac{dS_{r}}{dr}\right)^{2}=\frac{1}{\Delta}\left[(r^{2}+a^{2})E-aL_{z}\right]^{2}-\left[K+(L_{z}-aE)^{2}+\delta_{1}r^{2}\right]$ (11)
and
$\left(\frac{dS_{\theta}}{d\theta}\right)^{2}=K-\left(L_{z}^{2}\csc^{2}\theta-a^{2}E^{2}+\delta_{1}a^{2}\right)\cos^{2}\theta$ (12)
Here we abbreviate
$R=\left[(r^{2}+a^{2})E-aL_{z}\right]^{2}-\Delta\left[K+(L_{z}-aE)^{2}+\delta_{1}r^{2}\right]$ (13)
and
$\Theta=K-\left(L_{z}^{2}\csc^{2}\theta+a^{2}(\delta_{1}-E^{2})\right)\cos^{2}\theta$ (14)
Figure 1: The shadow of Kerr-scalaron black hole as observed by a distant observer at $\theta_{0}=\pi/2$ for spin 0.4 (left) and 0.99(right).Here KS denotes Kerr-scalaron.
The equations of motion are algebraically simplified to obtain the null geodesics for the variables. For null geodesics, $\delta_{1}=0$ and thus there is a possibility of minimizing the parameters to $\xi=L_{z}/E$ and $\eta=K/E^{2}$ which are often called as the impact parameters. Here $K$ denotes the Carter’s constant. The equations obtained are as follows:
$R=r^{4}+r^{2}(a^{2}-\xi^{2}-\eta)+2Mr[\eta+(\xi-a)^{2}]\left(1+\frac{1}{3}e^{-M_{\psi}r}\right)-a^{2}\eta$ (15)
and
$\Theta=\eta+a^{2}\cos^{2}\theta-\xi^{2}\cot^{2}\theta$ (16)
The equation of motion of the radial component of the null geodesics is represented by the function R(r) so that the circular photon orbits satisfies the conditions $R=0$ and $dR/dr=0$ (see, for example Chandrasekhar (1983)). We refrain from displaying the complex algebra involved here. The outcome of the calculations is the following set of impact parameters,
$\xi=\frac{M\alpha^{\prime}(r^{2}-a^{2})-r\Delta}{a(r-M\alpha^{\prime})}$ (17)
and
$\eta=\frac{r^{2}(4Ma^{2}\alpha^{\prime}+6Mr^{2}\alpha^{\prime}-9M^{2}r\alpha^{\prime}-r^{3})}{a^{2}(r-M\alpha^{\prime})^{2}}$ (18)
Here, $\alpha^{\prime}=1+\frac{1}{3}e^{-M_{\psi}r}$ and $\Delta=r^{2}+a^{2}-2Mr\alpha^{\prime}$. These two impact parameters define the celestial coordinates $\alpha$ and $\beta$ of the image as seen by the observer at infinity (Luminet (1979); Vázquez and Esteban (2004)) as
$\alpha=-\xi\csc\theta_{0}$ (19)
$\beta=\sqrt{\eta+a^{2}\csc^{2}\theta_{0}-\xi^{2}\cot^{2}\theta_{0}}$ (20)
Here $\theta_{0}$ is the inclination angle of the observer. The parametric curve ($(\alpha(\theta),\beta(\theta)$) delineates the shadow boundary as shown in Figure 1. The shadow morphology for the Kerr-scalaron metric exhibits dependence on the inclination angle of the observer and mass of the scalaron. It is very weakly dependent on spin except for the cases of extremely high spin as shown in the right panel of Figure 1. At extremely high spin, the presence of scalarons seems to suppress the spin induced distortion, with the shadow approaching the Kerr limit as the scalarons reach a mass of $\sim 10^{-16}$ eV.
2.2 Test of "no-hair" theorem
Testing the "no-hair" theorem requires measuring at least three different multipole moments. The size and shape of the black hole shadow are capable of measuring the mass, quadrupole and spin of the black hole and hence test the theorem observationally. We consider the parametrization of the Kerr quadrupole in terms of the parameter $\epsilon$ (called quadrupolar correction) as (Johannsen and Psaltis (2010))
$Q=-M(a^{2}+\epsilon M^{2})$ (21)
As $\epsilon\longrightarrow 0$ the quadrupole reduces to the Kerr value. For quantifying dependence of the shape of the shadow on the mass of the scalaron, inclination angle and deviation in the quadrupole moment, we observe the displacement and asymmetry of the black hole shadow. Figure 2 (a) shows the displacement as a function of the inclination angle for a fixed value of black hole spin. We follow the equation (Johannsen and Psaltis (2010))
$D=2a\sin i(1-0.41\epsilon\sin^{2}i)$ (22)
Here $D\equiv|x^{\prime}_{max}+x^{\prime}_{min}|/2$ is the horizontal displacement of the ring and $i$ is the inclination angle of the disk. The displacement is parameterized by the quadrupolar correction. The displacement is affected by scalaron mass which allows us to recognize a mapping between scalarons and quadrupolar correction. It increases with both increase in inclination angle and mass of the scalaron. A scalaron mass of $10^{-16}$ eV reproduces the displacement for quadrupolar correction $\epsilon=0.01$. This results in a value, Q = -0.17 which is only 1% away from the Kerr value for spin $a=0.4$.
The degree of asymmetry, quantified by the parameter $A/M$ is considered to be a direct test of the "no-hair" theorem (Johannsen and Psaltis (2010)). Figure 2 (b) illustrates the dependence of the asymmetry parameter on the inclination angle.
$\frac{A}{M}=\left[0.84\epsilon+0.36\left(\frac{a}{M}\right)^{3}\right]sin^{3/2}i$ (23)
Like displacement, the asymmetry is also mapped onto the scalaron mass. It may seem that the asymmetry is non negligible and has a high value for higher scalaron mass. However, the magnitude of asymmetry is less as compared to $0.36$ which is the Kerr limit (Johannsen and Psaltis (2010)). It is to be noted here, that while aligning the scalaron mass on displacement and asymmetry, we have respected the condition ($0.0\leq\epsilon\leq 0.5$) required for the parametrization of $D$ and $A/M$.
Mapping strong field bounds on gravity theory to the weak field bounds is performed by the PPN form of Schwarzschild limit of the Kerr-scalaron metric. The Eddington-Robertson-Schiff parameter $\gamma$ is related to the scalaron mass as (Amendola and Tsujikawa (2010)),
$\gamma=(3-e^{-M_{\psi}r})/(3+e^{-M_{\psi}r})$ (24)
Whereas low mass scalarons produce substantial departure from GR ($\gamma=1$), we see that the scalaron mass of $10^{-16}$ eV at the horizon scale ($\sim 0.1$ au) of Sgr A* reproduces the solar system constraint $\gamma-1=(2.1\pm 2.3)\times 10^{-5}$ given by the Cassini mission (Bertotti et al. (2003)) (see Figure 3). This confirms that $10^{-16}$ eV is a general relativistic mass limit for $f(R)$ gravity scalaron. This is close to earlier reports where it was shown that $10^{-16}$ eV scalarons are compatible with general relativistic periapsis shift of the S2 star encircling the GC black hole (Kalita (2020); Paul et al. (2023)). Scalarons of $10^{-17}-10^{-16}$ eV are also found to compatible with advance of perihelia of planets of our solar system (Paul and Kalita (2025))(see Figure 4 for studies converging to scalaron mass scale $10^{-16}$ eV). This exclusively demonstrates the possibility that the GR limit of the $f(R)$ theory is likely to be scale free.
(a) The displacement versus inclination angle for the Kerr-scalaron black hole for spin 0.4 and varying scalaron masses. The dashed lines denote the functional form used in Johannsen and Psaltis (2010) and the $\epsilon$ values are chosen so as to align the scalarons with predicted displacement.
(b) The asymmetry versus inclination angle for the Kerr-scalaron black hole for spin 0.4 and varying scalaron masses. The dashed lines denote the functional form used in Johannsen and Psaltis (2010) and the $\epsilon$ values are chosen so as to align the scalarons with predicted asymmetry.
Figure 2: Variation of displacement (a) and asymmetry (b) with inclination angle for a fixed value of black hole spin and various scalaron masses.
Figure 3: Variation of PPN parameter $\gamma$ with length scale for various scalaron masses.
Figure 4: Scalaron mass realized in other astrophysical tests, including planetary perihelia shift, EHT’s measurement of deviation parameter for shadow, S2’s periapsis shift and this study.
3 Scalaron mass from gravitational identifiers ($\kappa$, $\phi$)
Baker et al. (2015) identified the Kretschmann curvature ($\kappa$) and gravitational potential ($\phi$) environments where GR has been well tested and is yet to be tested with confidence. They used $\kappa$ and $\phi$ as gravitational identifiers which are related to each other by the relation(in appropriate units of $G$ and $c$)
$\kappa=\sqrt{48}\frac{c^{4}}{G^{2}M^{2}}\phi^{3}$ (25)
Here M is the GC black hole mass. The scalaron gravity is a natural consequence of curvature corrected vacuum fluctuations in horizon scale which predicts a novel relation between scalaron mass and black hole mass (Kalita (2018); Talukdar et al. (2024)),
$M_{\psi}=4.30\times 10^{-10}eV\left(\frac{M}{M_{\odot}}\right)^{-1}$ (26)
This allows us to relate scalaron mass with the gravitational identifiers as
$\left(\frac{M_{\psi}}{10^{-10}eV}\right)^{2}=48^{-1/2}\left(\frac{\kappa}{\phi^{3}}\right)\left(\frac{c^{4}}{G^{2}M_{\odot}^{2}}\right)^{-1}$ (27)
We consider the S-stars near the GC black hole which were considered by Baker et al. (2015) for estimating $\kappa$ and $\phi$ (see Table 3 of Baker et al. (2015)). We have used the range of $\kappa$ and $\phi$ experienced by the S-stars and find a unique value of $M_{\psi}=10^{-17}$ eV which is fairly close to the mass scale $M_{\psi}=10^{-16}$ predicted by Equation (22) and the GR limit compatible with black hole shadow size (Paul et al. (2024)) and shape (the present work). It is displayed in Figure 5. It inspires us to infer that the GR limit of $f(R)$ gravity produced in this work by shadow shape and hence "no-hair" theorem is consistent with the scalaron mass scale predicted by curvature and potential scales experienced by the S-stars. It has been found that orbits of S-stars and size and shape of the GC black hole are complimentary probes of $f(R)$ modified gravity.
We justify the mass scale $M_{\psi}\sim 10^{-17}-10^{-16}$ eV on the basis of an interesting observational bound on the scalaron field amplitude $\delta f^{\prime}(R_{o})=\psi_{o}-1$ where $\psi_{o}=f^{\prime}(R_{o})$, $\psi_{o}=1$ for GR with $f(R)=R$ and $R_{o}$ is the background Ricci curvature. Jain et al. (2013) used distance measurement in the local universe as a new test of gravity theories which obey chameleon screening mechanism (Khoury and Weltman (2004)). They used three distant indicators - cepheids, Tip of the Red Giant Branch (TRGB) and water masers which live in different gravitational environments ($\phi$), with modified gravity being screened upto different extents. The criterion for screening in a spherical mass distribution was given by Hu and Sawicki (2007) as $\phi>\frac{3}{2}\delta f^{\prime}(R_{o})$. It has been found that the background scalaron field values ($\delta f^{\prime}(R_{o})$) above $5\times 10^{-7}$ are ruled out at 95% confidence level. In this case $\psi_{o}=1$ to an incredible precision. Scalaron mass is related to $\psi_{o}$ and $\psi_{o}^{\prime}$ as (Amendola and Tsujikawa (2010))
$M_{\psi}^{2}=\frac{Ro}{3}\left(\frac{\psi_{o}}{\psi_{o}^{\prime}R_{o}}-1\right)$ (28)
If deviation from GR is small we can express the theory as
$f(R)=R+g(R)$ (29)
where $g(R)$ is slowly varying function of R. In exact general relativistic case $g(R)=-2\Lambda$ with $\Lambda$ being the cosmological constant. With this consideration, equation (28) gives
$\frac{\psi_{o}^{\prime}R_{o}}{\psi_{o}}\approx R_{o}g^{\prime\prime}(R_{o})<<1$ (30)
Therefore scalaron mass becomes
$M_{\psi}^{2}=\frac{1}{3}\left(\frac{\psi_{o}}{\psi_{o}^{\prime}}\right).$ (31)
With $\psi_{o}=1$ it gives $\psi_{o}^{\prime}=1/3M_{\psi}^{2}$. Since $\psi_{o}^{\prime}\approx\psi_{o}/R_{o}$ we get the following relation
$R_{o}\sim 3M_{\psi}^{2}\propto 1/M^{2}\propto\kappa$ (32)
This justifies the mass scale $10^{-17}-10^{-16}$ eV.
Figure 5: Variation of $M_{\psi}$ for identifiers $\kappa$ and $\phi$. The $\kappa$ and $\phi$ values for S-stars have been adopted from Table 3 of Baker et al. (2015).
4 Results and discussions
In this work we have investigated influence of scalaron on the shape of the black hole shadow and hence the black hole "no-hair" theorem. We employ the Kerr-scalaron metric to sketch the circular photon rings and their shapes. Keeping scalaron mass as a free parameter, we recognize quadrupolar correction which aligns with scalaron masses, reproducing displacement and asymmetry of the rings. The mass scale $10^{-16}$ eV at low black hole spin ($a=0.4$) is found to be compatible with Kerr black hole quadrupole upto 99.38% for both displacement and asymmetry. The maximum asymmetry is still far below the asymmetry allowed by the Kerr hypothesis, $A/M\sim 0.36$. This ensures validity of black hole "no-hair" theorem in presence of scalaron. The analysis of the gravitational identifiers have led us to infer that the scalaron mass scale predicted by these identifiers in the scale of S-stars is compatible with the one produced by black hole shadow analysis of shape and hence the "no-hair" theorem. We identify $10^{-16}$ eV scalaron mass as an appropriate general relativistic limit of $f(R)$ theory of gravity. Therefore the orbits of S-stars and the shape of the GC black hole shadow are complimentary probes of $f(R)$ theory of gravity. Scalaron mass of $10^{-16}$ eV is also found to be consistent with the earlier report (Paul et al. (2024)) of compatibility with size of the black hole shadow and the deviation parameter measured by the EHT.
The horizon scale of a black hole provides appropriate general relativistic limit of scalaron mass. The Yukawa correction ($1/3e^{-M_{\psi}r/r}$) to Newtonian potential at the shadow scale is only 2%. It further reduces after considering the upper bound on the Yukawa coupling strength, $g$ given by Abd El Dayem et al. (2025). It confirms the validity of Birkhoff’s theorem in the black hole environment - that the spacetime metric exterior to a spherically symmetric and static body has to be the Schwarzschild (zero spin limit). The scalaron field is massive. $10^{-16}$ eV corresponds to a Compton wavelength of $0.08$ au, the horizon size of the GC black hole. Gravity gets enhanced below this scale, manifesting as Yukawa fifth force. This clearly ensures validity of GR and GR limit of $f(R)$ gravity theory in astronomical scales such as S2’s orbit and black hole shadow. However, it keeps open the opportunity for gravitational wave observations involving merging of supermassive black holes which can be employed in near future to test scalaron gravity in a new regime. In the vicinity of compact objects, the coupling between gravitational waves and matter and modified dispersion relation of the gravitational waves due to massive gravity modes are two probes for strong field nature of gravity. The deviation of GR from modified theories of gravity becomes appreciable at low frequencies due to the presence of the massive modes of the theory. The deviation of results as compared to GR are non negligible for compact objects acting as lenses for gravitational waves. For example, LISA’s operating frequency band $\sim$ $10^{-4}$ Hz - $1$ Hz is eligible to detect $10^{-19}-10^{-15}$ eV scalarons through lensing of gravitational wave by $10^{3}-10^{6}M_{\odot}$ compact objects (Sharma et al. (2024)).
Solar system bound on the PPN parameter $\gamma$ is successfully reproduced by $10^{-16}$ eV scalarons. This scalaron mass is tantalizingly close to the one which is compatible with general relativistic periapsis shift of S2 star, size of the GC black hole shadow and advance of perihelia of planets in the solar system. The scalaron mass scale reported here is also compatible with the mass scale of the Yukawa fifth force mediator and ultralight dark matter ($10^{-21}-10^{-15}$ eV) coming from asteroid astrometric data and $10^{-18}-10^{-17}$ eV coming from analysis of the orbit of the asteroid Bennu (Tsai et al. (2023, 2024)). We conclude with an exciting possibility that general relativistic limit of $f(R)$ gravity theory is scale invariant. The scalaron mass scale presented here is likely to give us a new understanding of gravitational physics in a new regime of spacetime.
We wish to draw attention to few important future avenues of the results reported.
  • Existence of a GR limit of $f(R)$ scalaron gravity indicates that the theory can act as a close alternative to GR in cosmological setting.
  • Scalarons of the mass scale presented here can interestingly affect gravitational wave propagation and can be tested via future low frequency gravitational wave observatories.
[…]
  • The Compton wavelength of the scalarons ($\sim 0.08$ au) matches with the size of extra dimensions (5 dimensional Anti de-Sitter space) constrained by EHT measurements (Lemos et al. (2024)).The bound is $10^{-2}$ au. This is an important outcome for future avenues of the linkage between new gravitational physics and extra dimensions (see for example Nojiri and Odintsov (2003)).